### v7

Abstract
The efficient segmentation of text information from the background in degraded color document images is an important challenge in the preservation of ancient manuscripts. The imperfect preservation of ancient manuscripts has led to various types of degradation over time, such as staining, yellowing, and ink seepage, badly affecting document image binarization results. This work proposes a three-stage method to generate binarization image results for degraded colour document images using generative adversarial networks (GANs). Stage-1 involves applying discrete wavelet transform and retaining the low-low subband images for document image enhancement. In Stage-2, the original input image is split into red, green, and blue (RGB) three single-channel images and one grayscale image, and each image is trained with independent GANs to extract color foreground information. In Stage-3, the output images of Stage-2 and the resized input images are used to train independent GANs to generate document binarization results, enabling the combination of global and local features. The experimental results show that the Avg-Score of the proposed method is 77.64, 77.95, 79.05, 76.38, 75.34, and 77.00 on the (H)-DIBCO 2011, 2013, 2014, 2016, 2017, and 2018 datasets, which achieves the state-of-the-art level. The implementation code for this work is available at https://github.com/abcpp12383/ThreeStageBinarization.
keywords:
[…]
1 Introduction
Historical analyses of human societies primarily rely on reconstructing and analyzing preserved ancient documents. Parchment manuscripts and paintings have predominantly served as the means of preserving crucial ancient documents throughout the evolution of human society. Analyzing preserved ancient documents presents challenges due to potential forms of interference and degradation, including paper yellowing, text fading, page contamination and moisture damage.
Document image enhancement is one of the most important tasks in the reconstruction and analysis of ancient documents. It involves document image binarization to extract essential text information from the page. Elevating the image quality of deteriorated documents necessitates addressing various challenges, including text degradation and bleed-through issues, such as text degradation and bleed-through [1, 2, 3, 4]. However, traditional binarization methods cannot extract text information from the background of degraded documents effectively, often resulting in incomplete removal of shadows and noises, or loss of text information [5, 6, 7].
With the development of deep learning, neural networks have been applied to image binarization. Calvo-Zaragoza and Gallego [8] introduced a selectional auto-encoder approach, parsing the text through the trained network model, and performing binarization using threshold. Fully convolutional network (FCN) [9] enables the generated images to maintain the same size as the input images, and the task [10] of generating document binarization using FCN is also included in the semantic segmentation task. However, most image binarization methods primarily process grayscale documents, targeting contaminated black and white scanned ancient documents as their default.
Considering that the images of some ancient documents are in colour, we propose a novel three-stage method for document image enhancement and colour degradation document binarization. The proposed method consists of three stages. In Stage-1, we split the original input image into red, green, and blue (RGB) three single-channel images and one grayscale image, and apply discrete wavelet transform (DWT) to retain the low-low (LL) subband images and normalize them to achieve image enhancement. In Stage-2, we train four independent adversarial networks with the four single-channel images to eliminate background information and extract color foreground details from the local image patches. During the processing of Stage-3, the multiscale adversarial networks produce $224\times 224$ image patch (local) prediction outputs and $512\times 512$ resized image (global) prediction outputs for the document images, which are subsequently fused to obtain the final output.
To evaluate the performance of the proposed method for degraded document image binarization, we compare it with other methods using DIBCO datasets of 2011, 2013, 2014, 2016, 2017, and 2018 [11, 12, 13, 14, 15, 16].
The main contributions of this paper are as follows:
  • Introducing a novel three-stage generative adversarial networks method for enhancing and binarizing the degraded color documents.
  • The incorporation of DWT prior to the training of generative adversarial networks makes the image to be input closer to the Ground Truth image.
  • The proposed method achieves the state-of-the-art performance on multiple benchmark datasets.
The rest of this paper is organized as follows: Section 2 describes the deep learning method for document image binarization. Section 3 introduces a novel three-stage generative adversarial networks method. Section 4 presents the experiments performed on this work, and compares the performance of our method and other methods on several benchmark datasets. Finally, Section 5 discusses the conclusion and the future work of this paper.
2 Related Work
In previous researches focusing on document image binarization, the field is broadly divided into two categories: traditional image processing methods and deep learning-based semantic segmentation methods. However, traditional image processing binarization methods are always based on the pixel-level local threshold values, such as global binarization [6] and local adaptive binarization [5, 7]. These methods are not satisfactory when dealing with documents affected by excessive interference or degradation [17, 18].
With the advancement of deep learning, the convolutional neural network (CNN) framework [19, 20] has become the predominant backbone of computer vision tasks. Within computer vision, semantic segmentation is one of the most important research topics, encompassing document image binarization. For instance, Tensmeyer et al. [10] employed pseudo-F-measure (p-FM) and F-measure (FM) losses during the training of FCN to apply the model to the document image binarization task. Ronneberger et al. [21] extended FCN, presenting the U-Net architecture, which captures contextual and spatial information through its distinctive U-shaped design. Inspired by this, Peng et al. [22] introduced an encoder-decoder architecture for document image binarization, where the decoder transforms the low-resolution representation to the original dimensions, resulting in the binarized image. Vo et al. [23] first applied hierarchical deep supervised network (DSN) to document image binarization and achieved the state-of-the-art (SOTA) performance on several benchmark datasets. He et al. [24] proposed an iterative deep learning framework applied to document image binarization and the trained model achieved excellent performance.
In recent years, generative adversarial networks (GANs) [25] have achieved exciting results in the generation of real-world images. Unlike GANs which generate output images from random noise, Isola et al. [26] proposed Pix2Pix GAN based on conditional GAN (cGAN) [27] for image-to-image translation. To ensure that the generated images are more realistic, Pix2Pix GAN employs generators for loss function optimization. Within the network architecture, the generator evaluates the minimized generated image against the target output image, quantifying the $L1$ loss between them. Meanwhile, the discriminator assesses the credibility of the generated image, relying on the input image and the reference target image. This process is depicted by the subsequent equation:
$\begin{split}\emph{arg}\min_{G}\max_{D}V(G,D)=\mathop{\mathbb{E}_{x}}[\log(1-D(x,G(x)))]\\ +\mathop{\mathbb{E}_{x,y}}[\log D(x,y)]+\lambda\mathop{\mathbb{E}_{x,y}}[||y-G(x)||_{1}]\end{split}$ (1)
where the input data distribution $\mathop{\mathbb{P}_{x}}$ samples the input image $x$; the ground truth image of the input image is $y$, and the hyperparameter to increase the model regularization effect is $\lambda$. Generator $G$ generates the input image as the generated image $G(x)$. From Eq. (1), the loss function between the generated image and the ground truth image is $L1$ instead of $L2$, which is beneficial to reduce blurring and ambiguity in the generation process.
Bhunia et al. [28] performed document image enhancement on the DIBCO dataset and used cGAN [27] for image binarization. Inspired by Pix2Pix GAN [26], Zhao et al. [29] proposed a cascade generator structure for document image binarization based on Pix2Pix GAN to solve the problem of multi-scale information combination. De et al. [30] designed two discriminators to combine local and global information. Suh et al. [31] set up a two-stage generative adversarial network for image binarization using Patch GAN [26]. These methods keep breaking the SOTA performance on DIBCO datasets.
Figure 1: The architecture of Stage-1 of the proposed method. The original input image is divided into patches of $224\times 224$ size, and split into grayscale images and RGB three single-channel images. The single-channel images are all in grayscale. However, to visually represent the different channels of the image separately, the single-channel images in this diagram are replaced with images in three different RGB colors.
Figure 2: The architecture of the Haar wavelet transform used by the proposed method, which takes the red single-channel image patch as an example. The processes of the transform of the green and blue single-channel image patches are the same.
3 Proposed Method
3.1 Overall Network Architecture
Considering the complexity of color document image degradation and the advantages [31] of processing different color channels separately, our proposed method divides the overall network architecture into three stages, which focus on enhancing image quality of degraded color documents and extracting text information from the images. In Stage-1, we use DWT and normalization to preprocess the patches of the input image. These preprocessed image patches are employed to train as the input of the generator within the adversarial network framework. Stage-2 focuses on the elimination of background color noise and the extraction of color foreground information, while Stage-3 involves the fusion of local and global predictions to produce the final output. The detailed introduction of each stage can be found in Sections 3.2, 3.3, and 3.4, respectively.
3.2 Stage-1
This work employs four different generators to extract foreground color information from degraded document images while eliminating the background color. To preprocess the input of the different generators, we first split the original color image (24-bit) into RGB three single-channel images (8-bit) and one grayscale image (8-bit).
The ground truth (GT) image provided by the dataset is a single-channel binary image (8-bit), where pixel points have values of 0 for characters and 1 for the background. Images with varying channel numbers cannot be directly put into the discriminator for processing. As shown in Figure 1, we first split the RGB three channels of the original input image to create three single-channel RGB images. Then we apply DWT to each single-channel image, and retain the LL subband images for normalization. The GT images and the preprocessed grayscale images are summed at pixel wise as the corresponding ground truth images. Since the sizes of the retained LL subband images are $112\times 112$ (half of the $224\times 224$ length and width), we resize the images to the designated size of $224\times 224$ using bicubic interpolation [32].
3.2.1 Discrete Wavelet Transform
In this work, we use Haar wavelet transform [33], which is one of the DWT, to process the input image. This process can be conceived as a bandpass filter, selectively allowing signals with frequencies akin to those of the wavelet basis function to traverse. The flowchart of this DWT processing is illustrated in Figure 2, which generates four subband images: one low-frequency subband image and three high-frequency subband images. This work takes the single-channel split images as input and processes the n-direction by low-pass filtering, as shown in the following equation:
$\begin{split}v_{L}[m,n]=\sum_{k=0}^{K-1}x[m,2n-k]g[k]\end{split}$ (2)
Next, this work uses the low-pass filtering applied to the m-direction of Eq. (2) to obtain the following equation:
$\begin{split}v_{LL}[m,n]=\sum_{k=0}^{K-1}v_{L}[2m-k,n]g[k]\end{split}$ (3)
In contrast to the ground truth images provided by the dataset, the RGB single-channel images split from the original input images exhibit some substantial noise. This noise prevents them from being used directly in generating corresponding ground truth images through the generator and could potentially disrupt the performance of the trained adversarial network. DWT decomposes the image into both low-frequency and high-frequency information. The low-frequency component represents the mean value and the high-frequency component represents the difference value. The mean value, classified as low-frequency information, captures gradual changes and encodes both the picture’s contour information and approximate details. Conversely, the difference value, characterized as high-frequency information, captures rapid changes and encodes intricate picture details as well as local information (including noises). Therefore, we apply DWT and retain the LL subband images to input Stage-2, as shown in Figure 2, which effectively preserves contour information and reduces noise interferences.
Figure 3: Results of Haar wavelet transform & normalization of RGB single-channel images split from the example image patches.
Figure 4: Image normalization from image HW15 of DIBCO 2011 dataset, the above is an example of no normalization after DWT retains LL subband images, whereas the below is the example of normalization after DWT retains LL subband images.
3.2.2 Normalization
Normalization entails constraining data within a desired range, with the goal of simplifying the data processing and expediting convergence. The proposed equation for this process is expressed as:
$\begin{split}I_{N}=(\operatorname{IV_{nmax}}-\operatorname{IV_{nmin}})\frac{1}{1+e^{-\frac{I-\beta}{\alpha}}}+\operatorname{IV_{nmin}}\end{split}$ (4)
where IV is the intensity values, the image to be processed $I:\left\{\mathbb{X}\subseteq\mathbb{R}^{n}\right\}\rightarrow\{\operatorname{IV_{min}},\ldots,\operatorname{IV_{max}}\}$, and the generated image $I_{N}:\left\{\mathbb{X}\subseteq\mathbb{R}^{n}\right\}\rightarrow\{\operatorname{IV_{nmin}},\ldots,\operatorname{IV_{nmax}}\}$. $\alpha$ defines the width of the input intensity range, and $\beta$ defines the range centered intensity.
In the proposed method, we normalize the LL subband images to present initially incomparable data for comparison. The reason we choose the normalized LL subband image as the output is shown in Figure 3, where we apply DWT and normalization to each subband images.
The illustrative example of the impact of normalization on RGB single-channel images after the DWT operation is provided in Figure 4, when applied to image HW15 from the DIBCO 2011 dataset, normalization effectively balances the RGB. In contrast, without applying normalization results in an output image, it heavily skews towards the blue component. In addition, we also show that the global binarization results of the images using normalization are more similar to the Ground Truth images provided by the dataset compared to the images only processed by DWT, and the experiments are explained in detail in Section 4.5.
Figure 5: The architecture of Stage-2 of the proposed method. This method employs the U-Net++ [34, 35] architecture with EfficientNet-B6 [36] as the generator and improved PatchGAN [26] as the discriminator. The ground truth image during the training process is obtained from pixel-wise summation of the dataset provided image and the output from Stage-1.
Figure 6: The architecture of Stage-3 of the proposed method. This stage comprises both local and global predictions. The local prediction involves training several $224\times 224$ image patches, while the global prediction involves directly resizing the original image to $512\times 512$ for training.
3.3 Stage-2
In Stage-2, the proposed method is illustrated in Figure 5. Initially, the original input images are split into three RGB single-channel images and one grayscale image. Each of the single-channel images is input into an independent generator, resulting in the use of four generators in Stage-2 of the proposed method. These four generators train the model separately, enabling this method to remove background information from the localized image blocks while extracting colored foreground details. Additionally, all single-channel images share the same discriminator to distinguish between the generated image and its corresponding ground truth image.
3.3.1 Loss functions of GAN
To ensure a more robust convergence of the loss function, we follow the improved Wasserstein generative adversarial network with gradient penalty (WGAN-GP) [37] used by Suh et al. [31] applied to the objective function. Experiments by Bartusiak et al. [38] have demonstrated that the binary cross-entropy (BCE) loss yields superior performance compared to the $L1$ loss in binary classification tasks. Therefore, we chose the BCE loss instead of the $L1$ loss used in the Pix2Pix GAN [26]. The loss function of the WGAN-GP, which incorporates the BCE loss, is formulated as follows:
$\begin{split}\mathop{\mathbb{L}_{D}}=-\mathop{\mathbb{E}_{x,y}}[D(y,x)]+\mathop{\mathbb{E}_{x}}[D(G(x),x)]\\ +\alpha\mathop{\mathbb{E}_{x,\hat{y}\sim P_{\hat{y}}}}[(||\nabla_{\hat{y}}D(\hat{y},x)||_{2}-1)^{2}]\end{split}$ (5)
$\begin{split}\mathop{\mathbb{L}_{G}}=\mathop{\mathbb{E}_{x}}[D(G(x),x)]+\lambda\mathop{\mathbb{E}_{G(x),y}}[y\log G(x)\\ +(1-y)\log(1-G(x))]\end{split}$ (6)
where the penalty coefficient is $\alpha$, and the uniform sampling along a straight line between the ground truth distribution $P_{y}$ and the point pairs of the generated data distribution is $P_{\hat{y}}$. $\lambda$ is used to control the relative importance of different loss terms. In the discriminator, the generated image is distinguished from the real image by the loss function $\mathbb{L}_{D}$ in Eq. (5). In the generator, the distance between the generated image and the ground truth image in each color channel is minimized by the loss function $\mathbb{L}_{G}$ in Eq. (6).
3.3.2 Network architecture
Given the unpredictability of color document degradation, this work adopts four independent adversarial networks to extract text information from diverse colored backgrounds. This method mitigates color interference during document image binarization.
For the network proposed in Stage-2, four independent generators apply the same encoder-decoder architecture. The encoder extracts features, while a shared decoder is employed by multiple encoders of varying depths, and subsequently the features are concatenated. A skip connection is used between the encoder and the decoders. The encoder facilitates downsampling and context extraction, while the decoders undertake upsampling and amalgamate both upscaled and downsampled features. Specifically, this work employs EfficientNet-B6 [36] as the encoder, which is a lightweight CNN known for its performance in image classification tasks. For the discriminator, we choose the one [26, 29, 30, 39, 40, 41] used by Suh et al. [31] due to its superior generalization capabilities. Additionally, instead of U-Net [21] used by Suh et al. [31], this work uses its improved model U-Net++ [34, 35] for semantic segmentation, which can further improve the model performance.
The detailed procedure for training the proposed GAN models is illustrated in Algorithm 1. The training process is conducted in our training environment, which is described in Section 4.3, encompassing optimizer selection, learning rate configuration, and loss function definition. The algorithm demonstrates that the GAN model iteratively updates the loss function based on the authenticity assessment of the generated image by the discriminator. This iterative process enhances the ability of the generator to produce more real images.
Algorithm 1 Training process for the proposed GAN models
0:  Input images $I_{input}$
0:  $\omega=0.5$, $\alpha=10$, $\lambda=50$, $\eta=1\times 10^{-4}$
1:  Initialize: Parameters $\theta_{G_{r}}$, $\theta_{G_{g}}$, $\theta_{G_{b}}$, $\theta_{G_{gray}}$
2:  for number of training iterations do
3:     Split $I_{input}$ into $\{red,green,blue,gray\}$ images
4:     for $k=\{red,green,blue,gray\}$ do
5:        if $k$ is gray then
6:           $y_{k}\leftarrow y$
7:        else
8:           $x^{\prime}_{k}=Norm(V_{LL}(x_{k}))$
9:           $y_{k}\leftarrow x^{\prime}_{k}\bigcap y$
10:           Binarize $y_{k}$ with threshold value $t$.
11:        end if
12:        Update Discriminator $D$ (Eq. (5)):
13:        $\theta_{D}\leftarrow\theta_{D}-\eta_{D}\nabla_{\theta_{D}}\mathop{\mathbb{L}_{D}}$
14:        Update Generator $G_{k}$ (Eq. (6)):
15:        $\theta_{G_{k}}\leftarrow\theta_{G_{k}}-\eta_{G}\nabla_{\theta_{G_{k}}}\mathop{\mathbb{L}_{G}}$
16:     end for
17:  end for
18:  for $k=\{red,green,blue\}$ do
19:     $\hat{y}_{k}\leftarrow\omega G_{k}(x_{k})+(1-\omega)G_{gray}(x_{gray})$
20:  end for
21:  $\hat{y}\leftarrow[\hat{y}_{r},\hat{y}_{g},\hat{y}_{b}]$
3.4 Stage-3
In Stage-3, a multi-scale adversarial network is employed to generate images for both local and global binarization results, enhancing the distinction between the background and text with higher accuracy. Contrary to Stage-2, which predominantly relies on local prediction using the image patches ($224\times 224$), Stage-3 conducts global prediction on the resized input image ($512\times 512$). This method reduces the contextual information loss in text caused by the local predictions of Stage-2.
Figure 6 illustrates the process of two independent discriminators in Stage-3. Different from Stage-2 using 8-bit images for local prediction, Stage-3 uses 24-bit images as the input of the generator. Additionally, Figure 6 clarifies that the left-side input image comprises the image patches from the Stage-2 output, intended for local prediction. Conversely, the right-side input image represents the resized image from the original input for global prediction. In the binarization process of Stage-3, the network architecture of the generator remains consistent with Stage-2, differing only in the number of input channels.
4 Experiments
4.1 Datasets
4.1.1 DIBCO
The Document Image Binarization Competition (DIBCO) provides several datasets: DIBCO 2009 [42], H-DIBCO 2010 [43], DIBCO 2011 [11], H-DIBCO 2012 [44], DIBCO 2013 [12], H-DIBCO 2014 [13], H-DIBCO 2016 [14], DIBCO 2017 [15], and H-DIBCO 2018 [16]. These datasets comprise both grayscale and color images, including machine-printed as well as handwritten content. Participants in the competition were tasked with designing algorithms capable of extracting binarized text images from both handwritten and machine-printed images within these datasets.
4.1.2 PHIBD
The Persian Heritage Image Binarization Dataset (PHIBD) [45] contains 15 images of historical manuscripts from Mirza Mohammad Kazemaini’s old manuscript library in Yazd, Iran. The images within the dataset have suffered from various types of degradation.
4.1.3 SMADI
The Synchromedia Multispectral Ancient Document Images Dataset (SMADI) [1] consists of 240 multispectral images of 30 authentic historical handwritten letters. These letters, dating from the 17th to the 20th centuries, were written using iron gall ink. The original documents are deposited at the BAnQ Bibliotheque et Archives nationales du Quebec. A total of 240 images of real documents were meticulously captured, calibrated, and registered within this dataset. Each of these images was imaged with a CROMA CX MSI camera, resulting in 8 images per document.
4.1.4 Bickley Diary Dataset
The Bickley Diary Dataset (BD) [46] has been generously donated to the Singapore Methodist Archives by Erin Bickley. This dataset comprises diaries in which the images have been affected by both light and fold damages, thereby complicating the process of identification.
4.2 Evaluation Metric
In comparison with other traditional and SOTA methods for document image binarization, based on the case that DIBCO datasets do not provide OCR output, we use the following five evaluation metrics:
4.2.1 F-measure (FM)
${\rm FM}=\frac{2\times{\rm Recall}\times{\rm Precision}}{{\rm Recall}+{\rm Precision}}$ (7)
where the Recall is defined as $\frac{TP}{TP+FN}$; the Precision is defined as $\frac{TP}{TP+FP}$, and TP, FP, and FN represent the true positive, false positive, and false negative values, respectively.
4.2.2 Pseudo-F-measure (p-FM)
${\rm p-FM}=\frac{2\times{\rm Recall_{ps}}\times{\rm Precision_{ps}}}{{\rm Recall_{ps}}+{\rm Precision_{ps}}}$ (8)
where the Pseudo-Recall (${\rm Recall_{ps}}$) and the Pseudo-Precision (${\rm Precision_{ps}}$) both use the distance weights relative to the ground truth image profile to generate a new weighted ground truth image.
4.2.3 Peak signal-to-noise ratio (PSNR)
${\rm PSNR}=10\log(\frac{{\rm V}^{2}}{{\rm MSE}})$ (9)
where ${\rm V}$ represents the difference between the foreground and background, and the Mean Square Error (MSE) is defined as:
${\rm MSE}=\frac{\sum_{x=0}^{M-1}\sum_{y=0}^{N-1}{[}L(x,y)-L^{\prime}(x,y){]}^{2}}{M\times N}$ (10)
where $M$ and $N$ are the total number of pixels of image $L$ and image $L^{\prime}$ respectively. Generally, a higher PSNR value indicates greater similarity between the images.
4.2.4 Distance reciprocal distortion (DRD)
${\rm DRD}$ represents a measure of visual distortion in the binary image, which is shown in the following equation:
${\rm DRD}=\frac{{\rm\sum_{k}DRD_{k}}}{{\rm NUBN}}$ (11)
where ${\rm NUBN}$ corresponds to the number of uneven $8\times 8$ blocks in the ground truth image, and ${\rm DRD_{k}}$ is defined as:
${\rm DRD_{k}}=\sum_{i=-2}^{2}\sum_{j=-2}^{2}|{\rm G_{k}(i,j)-B_{k}(x,y)|\times N_{w}(i,j)}$ (12)
where ${\rm DRD_{k}}$ indicates the distortion of the ${\rm k_{th}}$ flipped pixel, calculated using a $5\times 5$ normalized weight matrix ${\rm N_{w}}$. ${\rm B_{k}}$ refers to the pixel of the binary image, and ${\rm G_{k}}$ denotes the pixel of the ground truth image.
4.2.5 Avg-Score (Avg)
The Peak Signal-to-Noise Ratio (PSNR) provides a straightforward and intuitive quantitative metric for comparing the generated image with the ground truth image. Although it can be used as a supplementary metric, it should not be the only evaluation measure. The F-measure and p-F-measure are used to assess the accuracy of different methods, but a higher value in these metrics does not always correspond to a higher PSNR value. In our experiments, it is common to observe that a method achieves the SOTA PSNR value, while the F-measure and p-F-measure values are lower than those of other models. To solve this problem, we followed the Avg-Score metric used by Jemni et al. [47] to compare the combined performance of different methods, as shown in the following equation:
$Avg=\frac{FM+p\!-\!FM+PSNR+(100-DRD)}{4}$ (13)
Table 1: Comparison of the PSNR values of the images from different training sets that have been processed using different methods with the corresponding ground truth images.
Method DIBCO 2009 H-DIBCO 2010 H-DIBCO 2012 Bickley Diary PHIBD SMADI Mean Values
Original Input 63.2879 61.3446 61.8542 57.6995 60.7355 61.2175 61.0232
DWT 64.1154 71.4422 64.8151 53.7037 51.5094 60.1374 60.9539
DWT&Norm 80.4286 86.1904 82.4625 74.2314 83.6403 78.5337 80.9145
Table 2: Quantitative comparison (FM/p-FM/PSNR/DRD/Avg) of the models trained with different patch sizes ($224\times 224$/$256\times 256$) on multiple benchmark datasets.
DIBCO 2011 DIBCO 2013 H-DIBCO 2014 H-DIBCO 2016 DIBCO 2017 DIBCO 2018 Mean Values
$256\times 256$ FM↑ 94.11 95.06 96.62 91.40 90.95 91.76 93.32
p-FM↑ 96.97 97.30 98.01 96.22 93.97 95.74 96.37
PSNR↑ 20.48 22.16 22.25 19.63 18.55 19.98 20.51
DRD↓ 1.76 1.66 0.96 2.96 2.95 2.75 2.17
Avg↑ 77.45 78.22 78.98 76.07 75.13 76.18 77.01
$224\times 224$ FM↑ 94.22 94.65 96.60 91.76 91.31 92.89 93.57
p-FM↑ 97.47 97.10 98.33 96.82 94.30 96.96 96.83
PSNR↑ 20.54 22.01 22.23 19.73 18.63 20.39 20.59
DRD↓ 1.69 1.98 0.97 2.81 2.88 2.23 2.09
Avg↑ 77.64 77.95 79.05 76.38 75.34 77.00 77.22
Figure 7: Visualization examples of the results of the proposed method at each stage. (a) the original input images, (b) the LL subband images obtained by Haar wavelet transform and normalization (Stage-1), (c) the enhanced images produced using our image enhancement method (Stage-2), (d) the binarized images obtained by combining local and global predictions (Stage-3), (e) the ground truth images.
Table 3: Comparison of performance evaluation of the proposed method and other methods for document binarization on multiple benchmark datasets. Best and 2nd best performance are in red
and blue colors, respectively. The data for the experimental results of other methods are obtained from [31, 48].
Method FM↑ p-FM↑ PSNR↑ DRD↓ Avg↑
Niblack [5] 70.44 73.03 12.39 24.95 57.73
Otsu [6] 78.60 80.53 15.31 22.57 62.97
Sauvola [7] 85.37 89.08 16.37 7.08 70.94
Howe [17] 94.04 95.06 20.43 2.10 76.86
Jia [18] 93.17 95.32 19.44 2.29 76.41
Vo [23] 92.58 94.67 19.16 2.38 76.01
He [24] 91.92 95.82 19.49 2.37 76.22
Suh [31] 93.57 95.93 20.22 1.99 76.93
Ours 94.22 97.47 20.54 1.69 77.64
(a) DIBCO 2011
Method FM↑ p-FM↑ PSNR↑ DRD↓ Avg↑
Niblack [5] 71.38 73.17 13.54 23.10 58.75
Otsu [6] 80.04 83.43 16.63 10.98 67.28
Sauvola [7] 82.71 87.74 17.02 7.64 69.96
Howe [17] 91.34 91.79 21.29 3.18 75.31
Jia[18] 93.28 96.58 20.76 2.01 77.15
Vo [23] 93.43 95.34 20.82 2.26 76.83
He [24] 93.36 96.70 20.88 2.15 77.20
Suh [31] 95.01 96.49 21.99 1.76 77.93
Ours 94.65 97.10 22.01 1.98 77.95
(b) DIBCO 2013
Method FM↑ p-FM↑ PSNR↑ DRD↓ Avg↑
Niblack [5] 86.01 88.04 16.54 8.26 70.58
Otsu [6] 91.62 95.69 18.72 2.65 75.85
Sauvola [7] 84.70 87.88 17.81 4.77 71.41
Howe [17] 96.49 97.38 22.24 1.08 78.76
Jia [18] 94.89 97.68 20.53 1.50 77.90
Vo [23] 95.97 97.42 21.49 1.09 78.45
He [24] 95.95 98.76 21.60 1.12 78.80
Suh [31] 96.36 97.87 21.96 1.07 78.78
Ours 96.60 98.33 22.23 0.97 79.05
(c) H-DIBCO 2014
Method FM↑ p-FM↑ PSNR↑ DRD↓ Avg↑
Niblack [5] 72.57 73.51 13.26 24.65 58.67
Otsu [6] 86.59 89.92 17.79 5.58 72.18
Sauvola [7] 84.64 88.39 17.09 6.27 70.96
Howe [17] 87.47 92.28 18.05 5.35 73.11
Jia [18] 90.10 93.72 19.00 4.03 74.70
Vo [23] 90.01 93.44 18.74 3.91 74.57
He [24] 91.19 95.74 19.51 3.02 75.86
Suh [31] 92.24 95.95 19.93 2.77 76.34
Ours 91.76 96.82 19.73 2.81 76.38
(d) H-DIBCO 2016
Method FM↑ p-FM↑ PSNR↑ DRD↓ Avg↑
Otsu [6] 77.73 77.89 13.85 15.54 63.48
Sauvola [7] 77.11 84.10 14.25 8.85 66.65
Howe [17] 90.10 90.95 18.52 5.12 73.61
Jia [18] 85.66 88.30 16.40 7.67 70.67
Suh [31] 90.95 94.65 18.40 2.93 75.27
Ours 91.31 94.30 18.63 2.88 75.34
(e) DIBCO 2017
Method FM↑ p-FM↑ PSNR↑ DRD↓ Avg↑
Otsu [6] 51.45 53.05 9.74 59.07 38.79
Sauvola [7] 67.81 74.08 13.78 17.69 59.50
Howe [17] 80.84 82.85 16.67 11.96 67.10
Jia [18] 76.05 80.36 16.90 8.13 66.30
Suh [31] 91.86 96.25 20.03 2.60 76.39
Ours 92.89 96.96 20.39 2.23 77.00
(f) DIBCO 2018
Figure 8: Example of document image binarization results from input image HW1 of DIBCO 2011 dataset by different methods: (a) original input, (b) Otsu [6], (c) Niblack [5], (d) Sauvola [7], (e) Vo [23], (f) He [24], (g) Suh [31], (h) Ours, (i) ground truth.
Figure 9: Example of document image binarization results from input image HW5 of DIBCO 2013 dataset by different methods: (a) original input, (b) Otsu [6], (c) Niblack [5], (d) Sauvola [7], (e) Vo [23], (f) He [24], (g) Suh [31], (h) Ours, (i) ground truth.
4.3 Implementation Details
4.3.1 Data preparation
The performance of a neural network depends on the quality and quantity of its training data. In general, the more training data enhances the better the performance of the neural network. To compare the performance of the proposed method with other methods, we follow the strategy described in [23, 24, 31] for constructing the training and test sets. Specifically, the training set includes a total of 143 images from DIBCO 2009 [42], H-DIBCO 2010 [43], H-DIBCO 2012 [44], PHIBD [45], SMADI [1], and BD [46] dataset. The test sets comprise a total of 82 images from DIBCO 2011 [11], DIBCO 2013 [12], H-DIBCO 2014 [13], H-DIBCO 2016 [14], DIBCO 2017 [15], and H-DIBCO 2018 [16].
To ensure a fair comparison of performance among all methods, we used the same data augmentation for our method and other SOTA methods. During the training process of the local prediction network, this work splits the input images into patches with $224\times 224$ in size. These patches were sampled with scale factors of ${0.75,1.25,1.5}$, subsequently resized to $224\times 224$, and rotated by 270°. In contrast, for training the global prediction network, this work employed scaling without rotation for data augmentation, using a scaling size of $512\times 512$. To diversify the training data, rotation augmentation was applied to the global prediction, involving rotations of 90°, 180°, and 270°, as well as horizontal and vertical flips.
4.3.2 Training
The experiments were conducted using the PyTorch framework in Python. We use NVIDIA RTX 3090 GPUs for all the following experiments.
In this study, we employ the generator comprising EfficientNet-B6 [36] with UNet++ [34, 35]. Regarding the training hyperparameters, Stage-2 and Stage-3 share similar settings, differing primarily in the number of training epochs: 150 epochs for global prediction and 10 epochs for the remaining networks. The choice of optimizer was Adam [49], with a learning rate set at $2\times 10^{-4}$. Additionally the generator employed $\beta_{1}=0.5$, while the discriminator used $\beta_{2}=0.999$.
4.3.3 Pre-training
To enhance training efficacy due to limited data availability, we employed EfficientNet-B6 [36] as the encoder in the generator, utilizing weights pretrained on the ImageNet dataset [50].
4.4 Comparison with other Methods
In our previous conference paper [51], we compared various deep learning-based methods and found that some of these methods [10, 28, 29] employed a “leave-one-out” strategy for selecting the training set. For instance, when the DIBCO 2013 dataset is used as the test set, all other datasets are utilized as the training set. Although our method still demonstrates superior performance at the SOTA level in such case [51], we consider it unfair to compare them directly because our model uses a smaller training set. Therefore, this work excludes these methods from this paper and only retains those that use the exact same training set for comparison.
4.5 Image Enhancement Experiment
In Stage-1, we applied DWT and normalization to the original input images. While mathematical theory [52] suggests that these processed images can retain contour information and reduce noise more efficiently, it is important to demonstrate their impact on experimental results. The previous conference paper [51] on this work presented ablation experiments on DWT and normalization. Based on this, it is better to intuitively show that, compared to the original input images, the DWT and normalization processed images are closer to the ground truth images.
In Stage-1 of this work, we set the threshold to 0.3 to binarize different single-channel images to obtain corresponding binarized images. In the experiment of this section, we applied the above image binarization to the original input images, the images processed with DWT, and the images processed with both DWT and normalization. Subsequently, we calculated the PSNR values of the binarized images compared to the ground truth images using these three methods, and the results are presented in Table 2. The PSNR mean values of the original input images are 61.0232. In contrast, the PSNR mean values of the images after DWT processing are slightly lower at 60.9539. However, the PSNR mean values of the images processed with both DWT and normalization are 80.9145, which means they are closer to the ground truth images. This demonstrates that it is more reasonable to choose the DWT and normalization processed image rather than the original input image and the DWT processed image.
4.6 Comparison Experiment
To ensure fairness in the experiments, we used the same training set and data augmentation techniques described in Section 4.3.1 for all the compared models. Unlike Suh et al. [31], who split the input images into $256\times 256$ patches, our proposed method uses $224\times 224$ patches. To demonstrate the positive impact of this change on the model performance, we conducted experiments comparing models using $224\times 224$ patches with those using $256\times 256$ patches, keeping all other settings unchanged.
Table 2 presents the performance of the models across different patch sizes on six DIBCO datasets. As established in Section 4.5, the images processed with DWT and normalization outperform those processed with either the original input images or only with DWT. Consequently, all models in this section utilize images processed with DWT and normalization. According to Table 2, the mean values of FM, p-FM, PSNR, DRD, and Avg for the model based on $224\times 224$ patches are 93.57, 96.83, 20.59, 2.09, and 77.22, respectively, which are superior to those for the model based on $256\times 256$ patches. This indicates that the model based on $224\times 224$ patches performs better, and therefore we will use this model for subsequent comparisons.
4.7 Experimental Results
Different from the conference paper [51] of this work, which uses option 3 to train the model, the proposed method in this paper applies DWT and normalization to the images in Stage-1 as input to train the model. The performance of the trained model will then be compared with that of other methods.
Figure 7 presents the visualization results at each stage of the trained model. Specifically, five images were randomly selected from the PHIBD and Bickley Diary datasets to demonstrate the step-by-step process of image enhancement and binarization using the proposed method. Figure 7 (b) represents the result of retaining the LL subband images after applying DWT and normalization. This stage performs noise reduction on the original input image, which is used as the ground truth image for the generator. Figure 7 (c) shows the result of image enhancement achieved using adversarial networks, where the image has the background color removed and the color of the text highlighted. Figure 7 (d) is the final output image obtained using the proposed whole method. Notably, it is evident that our final output result closely approximates Figure 7 (e), the ground truth image.
Table 3 presents the evaluation results of different models on six DIBCO datasets. Overall, the proposed method achieves the top two best performance in terms of FM, p-FM, PSNR, and DRD, with its Avg value reaching the SOTA level across all six datasets. Notably, the performance of the proposed method on the DIBCO 2011 and 2013 datasets is particularly good, obtaining the best values for each of the evaluated metrics. For instance, on the DIBCO 2011 dataset, the p-FM value of the proposed method is 97.47, significantly surpassing the second best value of 95.93. Similarly, on the DIBCO 2018 dataset, the FM value of the proposed method is 92.89, considerably higher than the second best value of 91.86. These results indicate that the proposed method is more effective than other methods for the binarization of degraded color documents.
In addition to assessing method performance by comparing quantitative metrics across various methods, this work aims to provide readers with an intuitive understanding of the distinctions by the visualization results. Consequently, we have chosen images from the DIBCO 2011 and 2013 datasets to illustrate examples of binarization results. Figures 8 and 9 show the binarization results of images HW1 and HW5 by different methods. These figures visually demonstrate that the proposed method outperforms deep learning based methods and the traditional binarization methods in terms of the shadow and noise elimination. Furthermore, the proposed method is good at preserving textual content while effectively mitigating the presence of the shadows and noise.
5 Conclusion
This paper introduces a novel three-stage method based on DWT and GANs to apply the task of text degradation in degraded document images with color backgrounds. The proposed method involves preserving the LL subband images through DWT to eliminate image noise, serving as the ground truth images for the generator. It trains independent networks utilizing RGB three single-channel images and one grayscale image, followed by a fusion of local and global prediction networks for binarization. As far as our knowledge extends, this work first represents the three-stage method that combines traditional image processing with deep learning for document binarization. The experimental results demonstrate that employing U-Net++ with EfficientNet-B6 as the generator and PatchGAN as the discriminator outperforms performance compared to other SOTA methods.
6 Declarations
This paper is an expanded paper from 20th Pacific Rim International Conference on Artificial Intelligence (PRICAI) held on November 15-19, 2023 in Jakarta, Indonesia.
6.1 Declaration of Funding
This research is supported by National Science and Technology Council of Taiwan, under Grant Number: NSTC 112-2221-E-032-037-MY2.
6.2 Declaration of competing interest
The authors have no financial or proprietary interests in any material discussed in this article.
6.3 Declaration of Generative AI and AI-assisted technologies in the writing process
The authors only use generative artificial intelligence (AI) and AI-assisted technologies to improve language.